Suppose That you are taking a course that has 4 tests. The first three tests are for 10 points each and the fourth test is for 200 points. to get a B in the course, you must have an average of at least 80 percent on the 4 tests. your scores on the first 3 tests were 76,72,84. what is the minimum score you need on the fourth test to get an B for the course?


Answer by checkley71(8405)

[(76+72+84)*.1+.7X=80
[232*.1+.7X]=80
(23.2+.7X)=80
23.2+.7X=80
.7X=80-23.2
.7X=56.8
X=56.8/.7
X=81.14 IS THE PERCENTAGE NEEDED ON THE FINAL TEST TO AVERAGE 80% ON ALL FOUR TESTS. .8114*200=162 IS REQUIRED ON THE FINAL 200 POINT TEST.
(76+72+84)/3*.1+.7*81.14=80
23.2+56=80
80=80